# Pure Message Passing Can Estimate Common Neighbor for Link Prediction

### v3

@@ 4.3 More scalable estimation
where $\tilde{{\bm{h}}}_{v}^{(l+1)}=\sum_{u\in\mathcal{N}_{v}}\tilde{{\bm{h}}}_{u}^{(l)}$ follows the message-passing defined in Equation 3. Such an estimation only requires the one-hop neighborhood $\mathcal{N}_{v}$, which is provided in a format of adjacency matrices/lists by most graph datasets. Therefore, we can substitute the structural features of MPLP with the estimation in Equation 7. We denote such a model with walk-level features as MPLP+.
Triangular substructure estimation
Our method, primarily designed to encapsulate the local structure of a target node pair, unexpectedly exhibits the capacity for estimating the count of triangles linked to individual nodes. This capability, traditionally considered beyond the reach of GNNs, marks a significant advancement in the field (Chen et al., 2020). Although triangle counting is less directly relevant in the context of link prediction, the implications of this capability are noteworthy. To maintain focus, we relegate the detailed discussion on pure message-passing for effective triangle counting to Appendix C.
@@ 5 Experiments
RA $68.25{\scriptstyle\pm 1.29}$ $72.29{\scriptstyle\pm 1.69}$ $28.05{\scriptstyle\pm 1.59}$ $40.77{\scriptstyle\pm 3.41}$ $64.00{\scriptstyle\pm 0.00}$ $\mathbf{{49.33}{\scriptstyle\pm{0.00}}}$ $51.98{\scriptstyle\pm 0.00}$
@@
ELPH $72.26{\scriptstyle\pm 2.58}$ $65.80{\scriptstyle\pm 2.26}$ $29.01{\scriptstyle\pm 2.66}$ $43.51{\scriptstyle\pm 2.37}$ $65.94{\scriptstyle\pm 0.58}$ OOM OOM
@@
Figure 4: Evaluation of model size and inference time on Collab. The inference time encompasses the entire cycle within a single epoch.
@@
We conduct evaluations across a diverse spectrum of 15 graph benchmark datasets, which include 8 non-attributed and 7 attributed graphs. It also includes three datasets from OGB (Hu et al., 2021) with predefined train/test splits. In the absence of predefined splits, links are partitioned into train, validation, and test sets using a 70-10-20 percent split. Our comparison spans three categories of link prediction models: (1) heuristic-based methods encompassing CN, AA, and RA; (2) node-level models like GCN and SAGE; and (3) link-level models, including SEAL, Neo-GNN (Yun et al., 2021), ELPH (Chamberlain et al., 2022), and NCNC (Wang et al., 2023). Each experiment is conducted 10 times, with the average score and standard deviations reported. The evaluation metrics are aligned with the standard metrics for OGB datasets, and we utilize Hits@50 for the remaining datasets. We limit the number of hops $r=2$, which results in a good balance of performance and efficiency. A comprehensive description of the experimental setup is available in Appendix D.
@@ Results
Model size and inference time
A separate assessment focuses on the trade-off between model size and inference time using the Collab dataset, with findings presented in Figure 4. Observing the prominent role of graph structure in link prediction performance on Collab, we introduce a streamlined version of our model, termed MPLP(no feat). This variant solely capitalizes on structural features, resulting in a compact model with merely 260 parameters. Nevertheless, its efficacy rivals that of models which are orders of magnitude larger. Furthermore, MPLP’s inference time for a single epoch ranks among the quickest in state-of-the-art approaches, underscoring its efficiency both in terms of time and memory footprint. More details can be found in Appendix D.3.
@@ 6 Conclusion
Broader Impact
@@ Appendix A Related work
Link prediction, inherent to graph data analysis, has witnessed a paradigm shift from its conventional heuristic-based methods to the contemporary, more sophisticated GNNs approaches. Initial explorations in this domain primarily revolve around heuristic methods such as CN, AA, RA, alongside seminal heuristics like the Katz Index (Katz, 1953), Jaccard Index (Salton & McGill, 1986), Page Rank (Brin & Page, 1998), and Preferential Attachment (Barabási & Albert, 1999). However, the emergence of graphs associated with node attributes has shifted the research landscape towards GNN-based methods. Specifically, these GNN-centric techniques bifurcate into node-level and link-level paradigms. Pioneers like Kipf & Welling introduce the Graph Auto-Encoder (GAE) to ascertain node pair similarity through GNN-generated node representation. On the other hand, link-level models, represented by SEAL (Zhang & Chen, 2018), opt for subgraph extractions centered on node pairs, even though this can present scalability challenges.
@@
The expressiveness of GNNs, particularly those of the MPNNs, has been the subject of rigorous exploration (Xu et al., 2018). A known limitation of MPNNs, their equivalence to the 1-Weisfeiler-Lehman test, often results in indistinguishable representation for non-isomorphic graphs. A suite of contributions has surfaced to boost GNN expressiveness, of which (Morris et al., 2021; Maron et al., 2020; Zhang & Li, 2021; Frasca et al., 2022) stand out. An elegant, yet effective paradigm involves symmetry-breaking through stochasticity injection (Sato et al., 2021; Abboud et al., 2021; Papp et al., 2021). Although enhancing expressiveness, such random perturbations can occasionally undermine generalizability. Diverging from these approaches, our methodology exploits probabilistic orthogonality within random vectors, culminating in a robust structural feature estimator that introduces minimal estimator variance.
@@
While node-level models like GAE offer enviable efficiency, they occasionally fall short in performance when compared with rudimentary heuristics (Chamberlain et al., 2022). Efforts to build scalable link-level alternatives have culminated in innovative methods such as Neo-GNN (Yun et al., 2021), which distills structural features from adjacency matrices for link prediction. Elsewhere, ELPH (Chamberlain et al., 2022) harnesses hashing mechanisms for structural feature representation, while NCNC (Wang et al., 2023) adeptly aggregates common neighbors’ node representation. Notably, DotHash (Nunes et al., 2023), which profoundly influenced our approach, employs quasi-orthogonal random vectors for set similarity computations, applying these in link prediction tasks.
Distinctively, our proposition builds upon, yet diversifies from, the frameworks of ELPH and DotHash. While resonating with ELPH’s architectural spirit, we utilize a streamlined, efficacious hashing technique over MinHash for set similarity computations. Moreover, we resolve ELPH’s limitations through strategic implementations like shortcut removal and norm rescaling. When paralleled with DotHash, our approach magnifies its potential, integrating it with GNNs for link predictions and extrapolating its applicability to multi-hop scenarios. It also judiciously optimizes variance induced by the structural feature estimator in sync with graph data. We further explore the potential of achieving higher expressiveness with linear computational complexity by estimating the substructure counting (Chen et al., 2020).
@@ Appendix C Estimate triangular substructures
Not only does MPLP encapsulate the local structure of the target node pair by assessing node counts based on varying shortest-path distances, but it also pioneers in estimating the count of triangles linked to any of the nodes— an ability traditionally deemed unattainable for GNNs (Chen et al., 2020). In this section, we discuss a straightforward implementation of the triangle estimation.
@@ C.1 Method
Constructing the structural feature with DE can provably enhance the expressiveness of the link prediction model (Li et al., 2020; Zhang et al., 2021). However, there are still prominent cases where labelling trick also fails to capture. Since labelling trick only considers the relationship between the neighbors and the target node pair, it can sometimes miss the subtleties of intra-neighbor relationships. For example, the nodes of $\text{DE}{(1,1)}$ in Figure 3 exhibit different local structures. Nevertheless, labelling trick like DE tends to treat them equally, which makes the model overlook the triangle substructure shown in the neighborhood. Chen et al. (2020) discusses the challenge of counting such a substructure with a pure message-passing framework. We next give an implementation of message-passing to approximate triangle counts linked to a target node pair—equivalent in complexity to conventional MPNNs.
@@ C.2 Experiments
Following the experiment in Section 6.1 of (Chen et al., 2020), we conduct an experiment to evaluate  MPLP’s ability to count triangular substructures. Similarly, we generate two synthetic graphs as the benchmarks: the Erdos-Renyi graphs and the random regular graphs. We also present the performance of baseline models reported in (Chen et al., 2020). Please refer to (Chen et al., 2020) for details about the experimental settings and baseline models. The results are shown in Table 3.
Table 3: Performance of different GNNs on learning the counts of triangles, measured by MSE divided by variance of the ground truth counts. Shown here are the median (i.e., third-best) performances of each model over five runs with different random seeds. Dataset Erdos-Renyi Random Regular GCN 8.27E-1 2.05 GIN 1.25E-1 4.74E-1 SAGE 1.48E-1 5.21E-1 sGNN 1.13E-1 4.43E-1 2-IGN 9.85E-1 5.96E-1 PPGN 2.51E-7 3.71E-5 LRP-1-3 2.49E-4 3.83E-4 Deep LRP-1-3 4.77E-5 5.16E-6 MPLP 1.61E-4 3.70E-4
Figure 5: Representation of the target link $(u,v)$ of MPLP after including the triangular estimation component. [Uncaptioned image]
@@ D.1 Benchmark datasets
  • USAir (Batagelj & Mrvar, 2006): a graph of US airlines;
  • NS (Newman, 2006): a collaboration network of network science researchers;
  • PB (Ackland & others, 2005): a graph of links between web pages on US political topics;
  • Yeast (Von Mering et al., 2002): a protein-protein interaction network in yeast;
  • C.ele (Watts & Strogatz, 1998): the neural network of Caenorhabditis elegans;
  • Power (Watts & Strogatz, 1998): the network of the western US’s electric grid;
  • Router (Spring et al., 2002): the Internet connection at the router-level;
  • E.coli (Zhang et al., 2018): the reaction network of metabolites in Escherichia coli.
4 out of 7 benchmarks with node attributes come from (Shchur et al., 2019), while Collab, PPA and Citation2 are from Open Graph Benchmark (Hu et al., 2021):
@@ D.2 More details in baseline methods
In our experiments, we explore advanced variants of the baseline models ELPH and NCNC. Specifically, for ELPH, Chamberlain et al. (2022) propose BUDDY, a link prediction method that preprocesses node representations to achieve better efficiency but compromises its expressiveness. NCNC (Wang et al., 2023) builds upon its predecessor, NCN, by first estimating the complete graph structure and then performing inference. In our experiments, we select the most expressiveness variant to make sure it is a fair comparison between different model architectures. Thus, we select ELPH over BUDDY, and NCNC over NCN to establish robust baselines in our study. We conduct a thorough hyperparameter tuning for ELPH and NCNC to select the best-performing models on each benchmark dataset. We follow the hyperparameter guideline of ELPH and NCNC to search for the optimal structures. For ELPH, we run through hyperparameters including dropout rates on different model components, learning rate, batch size, and dimension of node embedding. For NCNC, we experiment on dropout rates on different model components, learning rates on different model components, batch size, usage of jumping knowledge, type of encoders, and other model-specific terms like alpha. For Neo-GNN and SEAL, due to their relatively inferior efficiency, we only tune the common hyperparameters like learning rate, size of hidden dimensions.
@@ D.3 Evaluation Details: Inference Time
In Figure 4, we assess the inference time across different models on the Collab dataset for a single epoch of test links. Specifically, we clock the wall time taken by models to score the complete test set. This encompasses preprocessing, message-passing, and the actual prediction. For the SEAL model, we employ a dynamic subgraph generator during the preprocessing phase, which dynamically computes the subgraph. Meanwhile, for both ELPH and our proposed method, MPLP, we initially propagate the node features and signatures just once at the onset of inference. These are then cached for subsequent scoring sessions.
@@ D.4 Software and hardware details
We implement MPLP in Pytorch Geometric framework (Fey & Lenssen, 2019). We run our experiments on a Linux system equipped with an NVIDIA A100 GPU with 80GB of memory.
@@ D.6 Hyperparameters
  • Batch Size ($B$): We vary the batch size depending on the graph type: For the 8 non-attributed graphs, we explore batch sizes within $[512,1024]$. For the 4 attributed graphs coming from (Shchur et al., 2019), we search within $[2048,4096]$. For OGB datasets, we use $32768$ for Collab and PPA, and $261424$ for Citation2.
@@ F.1 Node label estimation accuracy and time
[…]
In Figure 8, we assess the accuracy of node label count estimation. For ELPH, the node signature dimension corresponds to the number of MinHash permutations. We employ a default hyperparameter setting for Hyperloglog, with $p=8$, a configuration that has demonstrated its adequacy in (Chamberlain et al., 2022). For time efficiency evaluation, we initially propagate and cache node signatures, followed by performing the estimation.

### v4

@@ 4.3 More scalable estimation
where $\tilde{{\bm{h}}}_{v}^{(l+1)}=\sum_{u\in\mathcal{N}_{v}}\tilde{{\bm{h}}}_{u}^{(l)}$ follows the message-passing defined in Equation 3. Similar to common GNNs, such a message-passing only requires the one-hop neighborhood $\mathcal{N}_{v}$, which is provided in a format of adjacency matrices/lists by most graph datasets. Therefore, we can substitute the structural features of MPLP with the estimation in Equation 7. We denote such a model with walk-level features as MPLP+.
4.4 Triangular substructure estimation
Our method, primarily designed to encapsulate the local structure of a target node pair, unexpectedly exhibits the capacity for estimating the count of triangles linked to individual nodes. This capability, traditionally considered beyond the reach of GNNs, marks a significant advancement in the field [32]. Although triangle counting is less directly relevant in the context of link prediction, the implications of this capability are noteworthy. To maintain focus, we relegate the detailed discussion on pure message-passing for effective triangle counting to Appendix C.
@@ 5 Experiments
RA $68.25{\scriptstyle\pm 1.29}$ $72.29{\scriptstyle\pm 1.69}$ $28.05{\scriptstyle\pm 1.59}$ $40.77{\scriptstyle\pm 3.41}$ $64.00{\scriptstyle\pm 0.00}$ $49.33{\scriptstyle\pm 0.00}$ $51.98{\scriptstyle\pm 0.00}$
@@
ELPH [^1] On OGB dataset Collab, PPA, and Citation2, we report the performance of BUDDY [16], a more streamlined version of ELPH. For the rest of the datasets, we report the performance of ELPH, which is empirically better when the computation budget is allowed. $72.26{\scriptstyle\pm 2.58}$ $65.80{\scriptstyle\pm 2.26}$ $29.01{\scriptstyle\pm 2.66}$ $43.51{\scriptstyle\pm 2.37}$ $65.94{\scriptstyle\pm 0.58}$ $\mathbf{{49.85}{\scriptstyle\pm{0.20}}}$ $87.56{\scriptstyle\pm 0.11}$
@@
We conduct evaluations across a diverse spectrum of 15 graph benchmark datasets, which include 8 non-attributed and 7 attributed graphs  [^2] Our code is publicly available at https://github.com/Barcavin/efficient-node-labelling.. It also includes three datasets from OGB [10] with predefined train/test splits. In the absence of predefined splits, links are partitioned into train, validation, and test sets using a 70-10-20 percent split. Our comparison spans three categories of link prediction models: (1) heuristic-based methods encompassing CN, AA, and RA; (2) node-level models like GCN and SAGE; and (3) link-level models, including SEAL, Neo-GNN [25], ELPH [16], and NCNC [19]. Each experiment is conducted 10 times, with the average score and standard deviations reported. The evaluation metrics are aligned with the standard metrics for OGB datasets, and we utilize Hits@50 for the remaining datasets. We limit the number of hops $r=2$, which results in a good balance of performance and efficiency. A comprehensive description of the experimental setup is available in Appendix D.
Table 3: Link prediction results on OGB datasets under HeaRT [33]. The top three models are colored by First, Second, Third.
Models Collab PPA Citation2
MRR Hits@20 MRR Hits@20 MRR Hits@20
CN 4.20 16.46 25.70 68.25 17.11 41.73
AA 5.07 19.59 26.85 70.22 17.83 43.12
RA 6.29 24.29 28.34 71.50 17.79 43.34
GCN 6.09 22.48 26.94 68.38 19.98 51.72
SAGE 5.53 21.26 27.27 69.49 22.05 53.13
SEAL 6.43 21.57 29.71 76.77 20.60 48.62
Neo-GNN 5.23 21.03 21.68 64.81 16.12 43.17
BUDDY 5.67 23.35 27.70 71.50 19.17 47.81
NCNC 4.73 20.49 33.52 82.24 19.61 51.69
MPLP+ 6.79 25.10 41.40 84.88 23.11 55.51
Figure 4: Evaluation of inference time on large-scale OGB datasets. The inference time encompasses the entire cycle within a full-batch inference.
@@ Results
We extend our evaluation of MPLP+ to assess its performance on large-scale datasets under the challenging HeaRT setting proposed by Li et al. [33]. HeaRT introduces a more rigorous and realistic set of negative samples during evaluation, typically resulting in a notable decline in performance across link prediction methods. As detailed in Table 3, MPLP+ consistently outperform all other methods across three OGB graph benchmarks in this demanding context. This underscores the robustness of MPLP+, affirming its ability to maintain superior performance across a variety of graph benchmarks and evaluation settings.
Time efficiency
We conduct an analysis of the time efficiency of our methods, MPLP and MPLP+, against established baselines using three large-scale OGB datasets. The results, illustrated in Figure 4, demonstrate that our approaches not only deliver superior performance across the graph benchmarks but also set a new benchmark for state-of-the-art time efficiency in full-batch inference. In particular, the primary component underlying our methods is the message-passing operation, which allows their inference speeds to rival that of the baseline GCN. Additionally, the structural feature estimations enhance the models’ expressiveness, enabling more accurate representation of graph structures, particularly in the context of link prediction tasks. More details can be found in Appendix D.3.
[…]
@@ 6 Conclusion
7 Acknowledgements
We would like to thank the anonymous reviewers for their insightful comments and helpful discussions. This research was supported in part by the University of Notre Dame’s Lucy Family Institute for Data and Society and the NSF Grant. We also acknowledge the Center for Research Computing at the University of Notre Dame for providing the computational resources that facilitated this study.
@@ Appendix A Related work
Link prediction, inherent to graph data analysis, has witnessed a paradigm shift from its conventional heuristic-based methods to the contemporary, more sophisticated GNNs approaches. Initial explorations in this domain primarily revolve around heuristic methods such as CN, AA, RA, alongside seminal heuristics like the Katz Index [34], Jaccard Index [35], Page Rank [36], and Preferential Attachment [24]. However, the emergence of graphs associated with node attributes has shifted the research landscape towards GNN-based methods. Specifically, these GNN-centric techniques bifurcate into node-level and link-level paradigms. Pioneers like Kipf and Welling introduce the Graph Auto-Encoder (GAE) to ascertain node pair similarity through GNN-generated node representation. On the other hand, link-level models, represented by SEAL [18], opt for subgraph extractions centered on node pairs, even though this can present scalability challenges.
@@
The expressiveness of GNNs, particularly those of the MPNNs, has been the subject of rigorous exploration [9]. A known limitation of MPNNs, their equivalence to the 1-Weisfeiler-Lehman test, often results in indistinguishable representation for non-isomorphic graphs. A suite of contributions has surfaced to boost GNN expressiveness, of which [37, 38, 39, 40] stand out. An elegant, yet effective paradigm involves symmetry-breaking through stochasticity injection [41, 42, 43]. Although enhancing expressiveness, such random perturbations can occasionally undermine generalizability. Diverging from these approaches, our methodology exploits probabilistic orthogonality within random vectors, culminating in a robust structural feature estimator that introduces minimal estimator variance.
@@
While node-level models like GAE offer enviable efficiency, they occasionally fall short in performance when compared with rudimentary heuristics [16]. Efforts to build scalable link-level alternatives have culminated in innovative methods such as Neo-GNN [25], which distills structural features from adjacency matrices for link prediction. Elsewhere, ELPH [16] harnesses hashing mechanisms for structural feature representation, while NCNC [19] adeptly aggregates common neighbors’ node representation. Notably, DotHash [12], which profoundly influenced our approach, employs quasi-orthogonal random vectors for set similarity computations, applying these in link prediction tasks.
Distinctively, our proposition builds upon, yet diversifies from, the frameworks of ELPH and DotHash. While resonating with ELPH’s architectural spirit, we utilize a streamlined, efficacious hashing technique over MinHash for set similarity computations. Moreover, we resolve ELPH’s limitations through strategic implementations like shortcut removal and norm rescaling. When paralleled with DotHash, our approach magnifies its potential, integrating it with GNNs for link predictions and extrapolating its applicability to multi-hop scenarios. It also judiciously optimizes variance induced by the structural feature estimator in sync with graph data. We further explore the potential of achieving higher expressiveness with linear computational complexity by estimating the substructure counting [32].
@@ Appendix C Estimate triangular substructures
Not only does MPLP encapsulate the local structure of the target node pair by assessing node counts based on varying shortest-path distances, but it also pioneers in estimating the count of triangles linked to any of the nodes— an ability traditionally deemed unattainable for GNNs [32]. In this section, we discuss a straightforward implementation of the triangle estimation.
@@ C.1 Method
Constructing the structural feature with DE can provably enhance the expressiveness of the link prediction model [26, 11]. However, there are still prominent cases where labelling trick also fails to capture. Since labelling trick only considers the relationship between the neighbors and the target node pair, it can sometimes miss the subtleties of intra-neighbor relationships. For example, the nodes of $\text{DE}{(1,1)}$ in Figure 3 exhibit different local structures. Nevertheless, labelling trick like DE tends to treat them equally, which makes the model overlook the triangle substructure shown in the neighborhood. Chen et al. [32] discusses the challenge of counting such a substructure with a pure message-passing framework. We next give an implementation of message-passing to approximate triangle counts linked to a target node pair—equivalent in complexity to conventional MPNNs.
@@ C.2 Experiments
Following the experiment in Section 6.1 of [32], we conduct an experiment to evaluate  MPLP’s ability to count triangular substructures. Similarly, we generate two synthetic graphs as the benchmarks: the Erdos-Renyi graphs and the random regular graphs. We also present the performance of baseline models reported in [32]. Please refer to [32] for details about the experimental settings and baseline models. The results are shown in Table 4.
Table 4: Performance of different GNNs on learning the counts of triangles, measured by MSE divided by variance of the ground truth counts. Shown here are the median (i.e., third-best) performances of each model over five runs with different random seeds.
Dataset Erdos-Renyi Random Regular
GCN 8.27E-1 2.05
GIN 1.25E-1 4.74E-1
SAGE 1.48E-1 5.21E-1
sGNN 1.13E-1 4.43E-1
2-IGN 9.85E-1 5.96E-1
PPGN 2.51E-7 3.71E-5
LRP-1-3 2.49E-4 3.83E-4
Deep LRP-1-3 4.77E-5 5.16E-6
MPLP 1.61E-4 3.70E-4
Figure 6: Representation of the target link $(u,v)$ of MPLP after including the triangular estimation component.
@@ D.1 Benchmark datasets
  • USAir [44]: a graph of US airlines;
  • NS [45]: a collaboration network of network science researchers;
  • PB [46]: a graph of links between web pages on US political topics;
  • Yeast [47]: a protein-protein interaction network in yeast;
  • C.ele [48]: the neural network of Caenorhabditis elegans;
  • Power [48]: the network of the western US’s electric grid;
  • Router [49]: the Internet connection at the router-level;
  • E.coli [50]: the reaction network of metabolites in Escherichia coli.
4 out of 7 benchmarks with node attributes come from [51], while Collab, PPA and Citation2 are from Open Graph Benchmark [10]:
@@ D.2 More details in baseline methods
In our experiments, we explore advanced variants of the baseline models ELPH and NCNC. Specifically, for ELPH, Chamberlain et al. [16] propose BUDDY, a link prediction method that preprocesses node representations to achieve better efficiency but compromises its expressiveness. NCNC [19] builds upon its predecessor, NCN, by first estimating the complete graph structure and then performing inference. In our experiments, we select the most expressiveness variant to make sure it is a fair comparison between different model architectures. Thus, we select ELPH over BUDDY, and NCNC over NCN to establish robust baselines in our study. We conduct a thorough hyperparameter tuning for ELPH and NCNC to select the best-performing models on each benchmark dataset. We follow the hyperparameter guideline of ELPH and NCNC to search for the optimal structures. For ELPH, we run through hyperparameters including dropout rates on different model components, learning rate, batch size, and dimension of node embedding. For NCNC, we experiment on dropout rates on different model components, learning rates on different model components, batch size, usage of jumping knowledge, type of encoders, and other model-specific terms like alpha. For Neo-GNN and SEAL, due to their relatively inferior efficiency, we only tune the common hyperparameters like learning rate, size of hidden dimensions.
@@ D.3 Evaluation Details: Inference Time
In Figure 4, we assess the inference time across different models on the OGB datasets for a single epoch of test links. Specifically, we clock the wall time taken by models to score the complete test set. This encompasses preprocessing, message-passing, and the actual prediction. For the SEAL model, we employ a dynamic subgraph generator during the preprocessing phase, which dynamically computes the subgraph. We substitute ELPH with BUDDY from [16] in this evaluation, since BUDDY exhibits better time efficiency compared to ELPH. For both BUDDY and our proposed methods, we initially propagate the node features and signatures just once at the onset of inference. These are then cached for subsequent scoring sessions.
@@ D.4 Software and hardware details
We implement MPLP in Pytorch Geometric framework [52]. We run our experiments on a Linux system equipped with an NVIDIA A100 GPU with 80GB of memory.
@@ D.6 Hyperparameters
  • Batch Size ($B$): We vary the batch size depending on the graph type: For the 8 non-attributed graphs, we explore batch sizes within $[512,1024]$. For the 4 attributed graphs coming from [51], we search within $[2048,4096]$. For OGB datasets, we use $32768$ for Collab and PPA, and $261424$ for Citation2.
@@ F.1 Node label estimation accuracy and time
In Figure 5, we assess the accuracy of node label count estimation. For ELPH, the node signature dimension corresponds to the number of MinHash permutations. We employ a default hyperparameter setting for Hyperloglog, with $p=8$, a configuration that has demonstrated its adequacy in [16]. For time efficiency evaluation, we initially propagate and cache node signatures, followed by performing the estimation.